# CC-Time: Cross-Model and Cross-Modality Time Series Forecasting

### v1

@@ 1 Introduction
With the rapid growth of the Internet of Things, vast amounts of time series data are being generated, driving increasing interest in time series forecasting (TSF) (Kaastra and Boyd 1996; Faloutsos et al. 2018). Current TSF methods primarily design specific modules to exploit the inherent knowledge of the time series data, and achieve good prediction accuracy (Liu et al. 2024c; Nie et al. 2023), which we call time-series-specific models in this paper.
@@
Recently, pre-trained language models (PLMs) have demonstrated remarkable success across diverse fields (Wang et al. 2024; Wu et al. 2024), prompting exploration in TSF (Zhou et al. 2023; Jin et al. 2024a). Some approaches attempt to leverage the representation capacity and sequential modeling capability of PLMs to capture time series patterns for TSF, which we call PLM-based models (Jin et al. 2024a; Liu et al. 2024d). Although these methods show good prospects, they have not yet achieved satisfactory prediction accuracy, leaving an under-explored problem of how to effectively activate the potential of PLMs for TSF. Motivated by this, we raise and explore two important questions:
@@
What time series characteristics could be modeled by pre-trained LMs? Real-world multivariate time series exhibit two critical characteristics: (1) temporal dependencies across time steps and (2) correlations with different channels. Capturing these features is essential for modeling the underlying data structure and improving prediction performance. However, existing PLM-based methods mainly focus on modeling temporal dependency and typically adopt a channel-independent approach, overlooking the potential of leveraging text modality knowledge stored in PLMs to model channel correlations (Zhou et al. 2023; Liu et al. 2024d). Meanwhile, time-series-specific methods are restricted to a single time-series modality, and they are more susceptible to numerical noise (Cheng et al. 2023) and lack the capacity to model correlations from other perspectives, such as the semantic perspective. Consequently, effectively modeling temporal dependency and channel correlations with multi-modality knowledge is necessary.
@@
Is relying solely on pre-trained LMs sufficient for building time series models? Recent studies indicate that PLM-based models and time-series-specific models focus on different aspects of modeling time series patterns(Jin et al. 2024a, b), and we also observe this phenomenon in the cross-model analysis of Section 4.3. Time-series-specific models excel at capturing basic patterns, such as trend and seasonality patterns (Jin et al. 2024b), and language models possess strong semantic understanding capability and multi-domain knowledge, which provide additional analytical perspectives for forecasting (Jin et al. 2024a; Liu et al. 2024d). Therefore, integrating these two types of models provides a more comprehensive understanding of time series patterns. However, a straightforward fusion, such as feature concatenation or knowledge distillation (Phuong and Lampert 2019; Kim and Rush 2016), does not bridge the gap between semantic information and numerical representations. Furthermore, due to the heterogeneity of time-series-specific models and PLMs, the correspondence between knowledge from both models is unclear. Therefore, a novel fusion method to adaptively integrate knowledge from both models is necessary.
@@ 2 Related Work
Channel correlation modeling has been proven to be essential for time series forecasting. Some existing methods adopt a channel-independent strategy (Nie et al. 2023; Lin et al. 2024; Xu, Zeng, and Xu 2024; Liu et al. 2024e), where the same weights are shared across all channels. Numerous studies employ Graph Neural Network (GNN) to capture the channel correlations (Liu et al. 2022; Yi et al. 2023; Shang, Chen, and Bi 2021). MTGNN (Wu et al. 2020) extends the application of GNN from spatio-temporal prediction to multivariate time series forecasting and proposes a method for computing an adaptive cross-channel graph. In addition to GNN-based methods, transformer-based models have made various attempts to capture correlations between channels (Zhang and Yan 2023; Liu et al. 2024c; Yang, Zhu, and Chen 2024; Wang et al. 2023). Crossformer (Zhang and Yan 2023) proposed a two-stage attention layer to capture the cross-time and cross-channel dependency efficiently. However, these models primarily design specific modules and only rely on time series modality to model correlations, which limits their capacity to fully capture complex channel correlations. CC-Time incorporates channel text descriptions and leverages language models to model channel correlations from a semantic perspective, empowering CC-Time’s completeness of modeling correlations.
@@
Pre-trained language models (PLMs) make progress in various fields beyond natural language processing. Recently, numerous studies have utilized powerful sequence modeling and representation capabilities of PLMs to model complex time series patterns, showcasing their potential in forecasting (Zhang et al. 2024; Jiang et al. 2024; Hu et al. 2025). These studies primarily involve direct usage (Gruver et al. 2023; Xue and Salim 2024), parameter-efficient fine-tuning (Zhou et al. 2023; Chang, Peng, and Chen 2023; Tan et al. 2024), prompting (Liu et al. 2024d, b; Cao et al. 2024; Pan et al. 2024), and modal alignment (Jin et al. 2024a; Sun et al. 2024; Liu et al. 2024a, 2025c, 2025b, 2025a). For example, GPT4TS (Zhou et al. 2023) fine-tunes the limited parameters of PLMs, demonstrating competitive performance by transferring knowledge from large-scale pre-training text data. UniTime (Liu et al. 2024b) designs domain instructions to align time series and text modality. Time-LLM (Jin et al. 2024a) reprograms time series into text to align the representation of PLMs. However, existing PLM-based methods primarily use PLMs as simple feature extractors and have not fully exploited their potential for modeling time series patterns. CC-Time proposes cross-model fusion to combine the strengths of both language models and time-series-specific models, adaptively integrating their knowledge to achieve a more holistic understanding of time series.
@@ 3.2 Cross-Modality Modeling with PLMs
In the PLM branch, we propose cross-modality modeling with PLMs to fully utilize multi-modality knowledge. Unlike existing PLM-based methods, we not only adopt a patching strategy to model temporal dependencies (Zhou et al. 2023), but also, for the first time, exploit PLMs to capture complex channel correlations from a semantic perspective. Specifically, as shown in the PLM branch of Figure 1, we construct a cross-modality embedding by integrating channel text descriptions with time series, providing semantic context for channel modeling. Based on the embedding, we design the PLM Channel Layer that models sample-specific channel correlations from a semantic perspective. Together with the global correlations captured by the Correlation Extractor from the entire training data, they jointly form comprehensive channel representations. Finally, the channel representations are fused with temporal representations extracted by the PLM Temporal Layer, yielding the final output representation at each PLM layer.
@@
Our proposed cross-modality embedding uses two modalities as the inputs: time series $\mathbf{X}_{\mathrm{{time}}}\in\mathbb{R}^{C\times T}$ and text descriptions $\mathbf{X}_{\mathrm{{text}}}\in\mathbb{R}^{C\times L}$, where $C$ represents the number of channels, and $T$ and $L$ represent the lengths of the time series and their text descriptions. For the input time series $\mathbf{X}_{\mathrm{{time}}}\in\mathbb{R}^{C\times T}$, channel embedding is used to describe the overall temporal properties of each channel, with a linear mapping along the temporal dimension yielding $\mathbf{E}_{\mathrm{{chan}}}\in\mathbb{R}^{C\times D_{l}}$, where $D_{l}$ represents the dimension of the embedded features. The corresponding text descriptions for each channel consist of two parts: channel semantic description and channel statistical information. The former provides a detailed semantic explanation for each channel, such as physical interpretations and causal relationships, and the latter offers relevant quantitative statistical details about these channels, such as the mean, variance. These descriptions about channels can be automatically acquired without requiring any additional human effort. The details about the channel text descriptions construction process are provided in Appendix A. To effectively integrate the embedding of the two modalities, we use linear mapping to compress the text embedding from the length $L$ to 1, aligning it with the dimension of channels. Then we add the text embedding and time series channels embedding to get the cross-modality embedding $\mathbf{E}_{\mathrm{{cross}}}\in\mathbb{R}^{C\times D_{l}}$.
@@
As illustrated in Figure 2, based on the cross-modality embedding as the input, the PLM channel layer leverages the PLMs’ knowledge to model sample-specific channel correlations from a semantic perspective. In parallel, we introduce a correlation extractor to capture global correlations across the entire training data. These two modules complement each other, enabling more comprehensive and robust modeling of channel correlations.
Given that channel correlations dynamically change over time, relying solely on the current time series sample and corresponding text descriptions to capture sample-specific correlations is not enough. To address this, the correlation extractor is designed to preserve global channel features across the entire training data while learning latent global correlations. Specifically, as shown in Figure 2, we initialize a learnable correlation extractor $\mathbf{E}\in\mathbb{R}^{C\times D_{r}}$ to preserve global channel features, where $D_{r}$ denotes the dimension of feature space. We use a linear projection to map the cross-modality embedding $\mathbf{E}_{\mathrm{{cross}}}$ into the feature space of the correlation extractor $\mathbf{E}$ as current channel features $\mathbf{E}_{\mathrm{f}}$.
Figure 2: Cross-modality correlation modeling with PLMs.
Subsequently, by utilizing the current channel features with the correlation extractor, a global correlation map $\mathbf{M}_{\mathrm{g}}\in\mathbb{R}^{C\times C}$ is adaptively learned. The process is as follows:
@@
where $\mathbf{M}_{\mathrm{g}}^{i,j}$ represents the weight of correlation between the $i$-th channel and the $j$-th channel. To make the correlations extracted from the extractor more global and generalizable, we use the current time series sample to update the correlation extractor. Specifically, we perform matrix multiplication between the global correlation map $\mathbf{M}_{\mathrm{g}}$ and the extractor $\mathbf{E}$ to get new channel features, followed by weighted integration with the existing channel features to produce the updated extractor $\mathbf{E^{\prime}}\in\mathbb{R}^{C\times D_{r}}$.
@@
$\mathbf{M}_{\mathrm{c}}=$ $\mathrm{Softmax}(\mathbf{Q}_{\mathrm{{chan}}}\mathbf{K}_{\mathrm{{chan}}}^{T}/\sqrt{D_{l}}),$ (2)
$\mathbf{Attn}_{\mathrm{{chan}}}$ $=(\epsilon\mathbf{M}_{\mathrm{g}}+(1-\epsilon)\mathbf{M}_{\mathrm{c}})\mathbf{V}_{\mathrm{{chan}}}$
@@
For temporal modeling, considering the sequential modeling capability of PLMs, we simultaneously leverage the pre-trained LM layers to capture complex temporal patterns, denoted as PLM temporal layers. Specifically, we adopt a patching strategy with PLMs (Zhou et al. 2023), where the time series $\mathbf{X}_{\mathrm{{time}}}\in\mathbb{R}^{C\times T}$ is divided into $N_{p}$ patches based on a patch size $S$. These patches are then passed through patch embedding to obtain $\mathbf{E}_{\mathrm{{patch}}}\in\mathbb{R}^{C\times N_{p}\times D_{l}}$, which is subsequently fed into the PLM temporal layer to model temporal dependencies, resulting in the final output $\mathbf{Z}_{\mathrm{{tem}}}\in\mathbb{R}^{C\times N_{p}\times D_{l}}$.
@@ 3.3 Cross-Model Fusion
Considering the advantage of PLMs and time-series-specific models, time-series-specific models focus on modeling patterns from numerical representations, while PLMs demonstrate strong generalization and complex-patterns-modeling capabilities, which provide additional perspectives for forecasting. Motivated by this, we propose the cross-model fusion to adaptively integrate knowledge from the two types of models, thus fully leveraging their respective strengths. Our proposed fusion module performs a two-step fusion: 1) adaptively fusing hierarchical features among layers from the PLMs to enhance the richness of the PLM representations, and 2) cross-model fusion between the PLM branch and the time series branch to construct a comprehensive understanding of time series.
@@ 4 Experiment
We select eight state-of-the-art models for time series forecasting as baselines, including PLM-based models: S2IP-LLM (Pan et al. 2024), FSCA(Hu et al. 2025), Time-LLM (Jin et al. 2024a), UniTime (Liu et al. 2024b), and GPT4TS (Zhou et al. 2023), and time-series-specific models: iTransformer (Liu et al. 2024c), Crossformer (Zhang and Yan 2023), PatchTST (Nie et al. 2023).
Experiment Settings
For a fair comparison, we set the input length $T$ to 96 and the output length $F$ to 96, 192, 336, and 720 for all baseline models and CC-Time. At the same time, all models do not use the drop last strategy (Qiu et al. 2024). Refer to GPT4TS (Zhou et al. 2023), we set GPT2 as the default architecture for the PLM branch of CC-Time and all pre-trained LM-based baselines. Meanwhile, in the PLM Layers of CC-Time, we only fine-tune the positional encoding and layer normalization to reduce learnable parameters.
@@ 4.1 Main Results
Table 1: Time series forecasting results with the input length $T=96$ and the prediction length $F=\{96,192,336,720\}$. Bold: the best and underline: the second best. Complete results are in Table 10 of the Appendix, and comparisons with time series foundation models are provided in Appendix D.
@@
We conduct few-shot forecasting using 10% of the training data on the ETT and Weather datasets to assess the few-shot learning capabilities of PLMs. As shown in Table 2, our proposed CC-Time achieves state-of-the-art prediction accuracy. Overall, CC-Time and PLM-based methods significantly outperform time-series-specific methods like PatchTST and iTransformer. This indicates that PLM-based methods have strong generalization capabilities, making them well-suited for scenarios with limited or sparse time series data, highlighting the potential of PLMs for TSF.
@@
Table 2: 10% few shot forecasting results with the input length T = 96 and the prediction length F = 96, 192, 336, 720. Bold: the best and underline: the second best. Full results are in Table 11 of the Appendix.

### v3

@@ 1 Introduction
With the rapid growth of the Internet of Things, vast amounts of time series data are being generated, driving increasing interest in time series forecasting (TSF) Kaastra & Boyd (1996); Faloutsos et al. (2018). Current TSF methods primarily design specific modules to exploit the inherent knowledge of the time series data, and achieve good prediction accuracy Liu et al. (2024c); Nie et al. (2023), which we call time-series-specific models in this paper.
@@
Recently, pre-trained language models (PLMs) have demonstrated remarkable success across diverse fields Wang et al. (2024); Wu et al. (2024), prompting exploration in TSF Zhou et al. (2023); Jin et al. (2024a). Some approaches attempt to leverage the representation capacity and sequential modeling capability of PLMs to capture time series patterns for TSF, which we call PLM-based models Liu et al. (2024d). Although these methods show good prospects, they have not yet achieved satisfactory prediction accuracy, leaving an under-explored problem of how to effectively activate the potential of PLMs for TSF. Motivated by this, we raise and explore two important questions:
@@
What time series characteristics could be modeled by pre-trained LMs? Real-world multivariate time series exhibit two critical characteristics: (1) temporal dependencies across time steps and (2) correlations with different channels. Capturing these features is essential for modeling the underlying data structure and improving prediction performance. However, existing PLM-based methods mainly focus on modeling temporal dependency and typically adopt a channel-independent approach, overlooking the potential of leveraging text modality knowledge stored in PLMs to model channel correlations Zhou et al. (2023); Liu et al. (2024d). Meanwhile, time-series-specific methods are restricted to a single time-series modality, and they are more susceptible to numerical noise Cheng et al. (2023) and lack the capacity to model correlations from other perspectives, such as the semantic perspective. Consequently, effectively modeling temporal dependency and channel correlations with multi-modality knowledge is necessary.
@@
Is relying solely on pre-trained LMs sufficient for building time series models? Recent studies indicate that PLM-based models and time-series-specific models focus on different aspects of modeling time series patternsJin et al. (2024a; b), and we also observe this phenomenon in the cross-model analysis of Section 4.4. Time-series-specific models excel at capturing basic patterns, such as trend and seasonality patterns Jin et al. (2024b), and language models possess strong semantic understanding capability and multi-domain knowledge, which provide additional analytical perspectives for forecasting Jin et al. (2024a); Liu et al. (2024d). Therefore, integrating these two types of models provides a more comprehensive understanding of time series patterns. However, a straightforward fusion, such as feature concatenation or knowledge distillation Phuong & Lampert (2019); Kim & Rush (2016), does not bridge the gap between semantic information and numerical representations. Furthermore, due to the heterogeneity of time-series-specific models and PLMs, the correspondence between knowledge from both models is unclear. Therefore, a novel fusion method to adaptively integrate knowledge from both models is necessary.
@@ 2 Related Work
Channel correlation modeling has been proven to be essential for time series forecasting. Some existing methods adopt a channel-independent strategy Nie et al. (2023); Lin et al. (2024); Xu et al. (2024); Liu et al. (2024e), where the same weights are shared across all channels. Numerous studies employ Graph Neural Network (GNN) to capture the channel correlations Liu et al. (2022); Yi et al. (2023); Shang et al. (2021). MTGNN Wu et al. (2020) extends the application of GNN from spatio-temporal prediction to multivariate time series forecasting and proposes a method for computing an adaptive cross-channel graph. In addition to GNN-based methods, transformer-based models have made various attempts to capture correlations between channels Zhang & Yan (2023); Liu et al. (2024c); Yang et al. (2024); Wang et al. (2023). Crossformer Zhang & Yan (2023) proposed a two-stage attention layer to capture the cross-time and cross-channel dependency efficiently. However, these models primarily design specific modules and only rely on time series modality to model correlations, which limits their capacity to fully capture complex channel correlations. CC-Time incorporates channel text descriptions and leverages language models to model channel correlations from a semantic perspective, empowering CC-Time’s completeness of modeling correlations.
@@
Pre-trained language models (PLMs) make progress in various fields beyond natural language processing. Recently, numerous studies have utilized powerful sequence modeling and representation capabilities of PLMs to model complex time series patterns, showcasing their potential in forecasting Zhang et al. (2024); Jiang et al. (2024); Hu et al. (2025). These studies primarily involve direct usage Gruver et al. (2023); Xue & Salim (2024), parameter-efficient fine-tuning Zhou et al. (2023); Chang et al. (2023); Tan et al. (2024), prompting Liu et al. (2024d; b); Cao et al. (2024); Pan et al. (2024), and modal alignment Jin et al. (2024a); Sun et al. (2024); Liu et al. (2024a; 2025c; 2025b; 2025a). For example, GPT4TS Zhou et al. (2023) fine-tunes the limited parameters of PLMs, demonstrating competitive performance by transferring knowledge from large-scale pre-training text data. UniTime Liu et al. (2024b) designs domain instructions to align time series and text modality. Time-LLM Jin et al. (2024a) reprograms time series into text to align the representation of PLMs. However, existing PLM-based methods primarily use PLMs as simple feature extractors and have not fully exploited their potential for modeling time series patterns. CC-Time proposes cross-model fusion to combine the strengths of both language models and time-series-specific models, adaptively integrating their knowledge to achieve a more holistic understanding of time series.
@@ 3.2 Cross-Modality Modeling with PLMs
In the PLM branch, we propose cross-modality modeling with PLMs to fully utilize multi-modality knowledge. Unlike existing PLM-based methods, we not only adopt a patching strategy to model temporal dependencies Zhou et al. (2023), but also, for the first time, exploit PLMs to capture complex channel correlations from a semantic perspective. Specifically, as shown in the PLM branch of Figure 1, we construct a cross-modality embedding by integrating channel text descriptions with time series, providing semantic context for channel modeling. Based on the embedding, we design the PLM Channel Layer that models sample-specific channel correlations from a semantic perspective. Together with the global correlations captured by the Correlation Extractor from the entire training data, they jointly form comprehensive channel representations. Finally, the channel representations are fused with temporal representations extracted by the PLM Temporal Layer, yielding the final output representation at each PLM layer.
@@
Our proposed cross-modality embedding uses two modalities as the inputs: time series $\mathbf{X}_{\mathrm{{time}}}\in\mathbb{R}^{C\times T}$ and text descriptions $\mathbf{X}_{\mathrm{{text}}}\in\mathbb{R}^{C\times L}$, where $C$ represents the number of channels, and $T$ and $L$ represent the lengths of the time series and their text descriptions. For the input time series $\mathbf{X}_{\mathrm{{time}}}\in\mathbb{R}^{C\times T}$, channel embedding is used to describe the overall temporal properties of each channel, with a linear mapping along the temporal dimension yielding $\mathbf{E}_{\mathrm{{chan}}}\in\mathbb{R}^{C\times D_{l}}$, where $D_{l}$ represents the dimension of the embedded features. The corresponding text descriptions for each channel consist of two parts: channel semantic description and channel statistical information. The former provides a detailed semantic explanation for each channel, such as physical interpretations and causal relationships, and the latter offers relevant quantitative statistical details about these channels, such as the mean, variance. These descriptions about channels can be automatically acquired without requiring any additional human effort. The details about the channel text descriptions construction process are provided in Appendix A. To effectively integrate the embedding of the two modalities, we use linear mapping to compress the text embedding, aligning it with the dimension of channels. Then we add the text embedding and time series channels embedding to get the cross-modality embedding $\mathbf{E}_{\mathrm{{cross}}}\in\mathbb{R}^{C\times D_{l}}$.
@@
Figure 2: Cross-modality correlation modeling.
As illustrated in Figure 2, based on the cross-modality embedding as the input, the PLM channel layer leverages the PLMs’ knowledge to model sample-specific channel correlations from a semantic perspective. In parallel, we introduce a correlation extractor to capture global correlations across the entire training data. These two modules complement each other, enabling more comprehensive and robust modeling of channel correlations. Given that channel correlations dynamically change over time, relying solely on the current time series sample and corresponding text descriptions to capture sample-specific correlations is not enough. To address this, the correlation extractor is designed to preserve global channel features across the entire training data while learning latent global correlations. Specifically, as shown in Figure 2, we initialize a learnable correlation extractor $\mathbf{E}\in\mathbb{R}^{C\times D_{r}}$ to preserve global channel features, where $D_{r}$ denotes the dimension of feature space. We use a linear projection to map the cross-modality embedding $\mathbf{E}_{\mathrm{{cross}}}$ into the feature space of the correlation extractor $\mathbf{E}$ as current channel features $\mathbf{E}_{\mathrm{f}}$. Subsequently, by utilizing the current channel features with the correlation extractor, a global correlation map $\mathbf{M}_{\mathrm{g}}\in\mathbb{R}^{C\times C}$ is adaptively learned. The process is as follows:
@@
where $\mathbf{M}_{\mathrm{g}}^{i,j}$ represents the weight of correlation between the $i$-th channel and the $j$-th channel. To make the correlations extracted from the extractor more global and generalizable, we use the current time series sample to update the correlation extractor. Specifically, we perform matrix multiplication between the global correlation map $\mathbf{M}_{\mathrm{g}}$ and the extractor $\mathbf{E}$ to get new channel features, followed by weighted integration with the existing channel features to produce the updated extractor $\mathbf{E^{\prime}}$.
@@
$\mathbf{M}_{\mathrm{c}}=$ $\mathrm{Softmax}(\mathbf{Q}_{\mathrm{{chan}}}\mathbf{K}_{\mathrm{{chan}}}^{T}/\sqrt{D_{l}}),\quad\mathbf{Attn}_{\mathrm{{chan}}}$ $=(\epsilon\mathbf{M}_{\mathrm{g}}+(1-\epsilon)\mathbf{M}_{\mathrm{c}})\mathbf{V}_{\mathrm{{chan}}}$ (2)
@@
For temporal modeling, considering the sequential modeling capability of PLMs, we simultaneously leverage the pre-trained LM layers to capture complex temporal patterns, denoted as PLM temporal layers. Specifically, we adopt a patching strategy with PLMs Zhou et al. (2023), where the time series $\mathbf{X}_{\mathrm{{time}}}\in\mathbb{R}^{C\times T}$ is divided into $N_{p}$ patches based on a patch size $S$. These patches are then passed through patch embedding to obtain $\mathbf{E}_{\mathrm{{patch}}}\in\mathbb{R}^{C\times N_{p}\times D_{l}}$, which is subsequently fed into the PLM temporal layer to model temporal dependencies, resulting in the final output $\mathbf{Z}_{\mathrm{{tem}}}\in\mathbb{R}^{C\times N_{p}\times D_{l}}$.
@@ 3.3 Cross-Model Fusion
Considering the advantage of PLMs and time-series-specific models, specific models focus on modeling patterns from numerical representations, while PLMs demonstrate strong generalization and complex-patterns-modeling capabilities, which provide additional perspectives for forecasting. Motivated by this, we propose the cross-model fusion to adaptively integrate knowledge from the two types of models, thus fully leveraging their respective strengths. Our proposed fusion module performs a two-step fusion: 1) adaptively fusing hierarchical features among layers from the PLMs to enhance the richness of the PLM representations, and 2) cross-model fusion between the PLM branch and the time series branch to construct a comprehensive understanding of time series.
@@ 4.1 Experimental Setup
We select eight state-of-the-art models for time series forecasting as baselines, including PLM-based models: S2IP-LLM Pan et al. (2024), FSCAHu et al. (2025), Time-LLM Jin et al. (2024a), UniTime Liu et al. (2024b), and GPT4TS Zhou et al. (2023), and time-series-specific models: iTransformer Liu et al. (2024c), Crossformer Zhang & Yan (2023), PatchTST Nie et al. (2023).
Settings
For a fair comparison, we set the input length $T$ to 96 and the output length $F$ to 96, 192, 336, and 720 for all baseline models and CC-Time. At the same time, all models do not use the drop last strategy Qiu et al. (2024). Refer to GPT4TS Zhou et al. (2023), we set GPT2 as the default architecture for the PLM branch of CC-Time and all pre-trained LM-based baselines. Meanwhile, in the PLM Layers of CC-Time, we only fine-tune the positional encoding and layer normalization to reduce learnable parameters.
@@ 4.2 Main Results
Table 1: Time series forecasting results with the input length $T=96$ and the prediction length $F=\{96,192,336,720\}$. Bold: the best and underline: the second best. Complete results are in Table 10, and comparisons with time series foundation models are provided in Appendix D.
@@
We conduct few-shot forecasting using 10% of the training data on the ETT and Weather datasets to assess the few-shot learning capabilities of PLMs. As shown in Table 2, our proposed CC-Time achieves state-of-the-art prediction accuracy. Overall, CC-Time and PLM-based methods significantly outperform time-series-specific methods like PatchTST and iTransformer. This indicates that PLM-based methods have strong generalization capabilities, making them well-suited for scenarios with limited or sparse time series data, highlighting the potential of PLMs for time series forecasting.
@@
Table 2: 10% few shot forecasting results with the input length T = 96 and the prediction length F = {96, 192, 336, 720}. Bold: the best and underline: the second best. Full results are in Table 11.